OpenStudy (anonymous):

Choose the correct graph of the given system of equations. y - 2x = -1 x + 3y = 4 graph of two lines, one with a positive slope and one with a negative slope, that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

we can gather the slope of these equations using -a/b so the first equation ahs a slope of 2/1 and the 2nd has a slope of -3/1

OpenStudy (anonymous):

so they are not parallel

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

now test x=-1 and x=1 to see if they give the y-value shown in the point

OpenStudy (anonymous):

but how do you find the y value

OpenStudy (anonymous):

plug x and and if the y value agrees with the ones in the points for A or B A (-1,1) when x=-1 y=1 B (1,1) when x=1 y=1

OpenStudy (anonymous):

plug x into both equations and see if you get the correct y value for both

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so lets plug x and -1 y - 2x = -1 x + 3y = 4 y-2(-1)=-1 -1+3y=4 y+2=-1 3y=5 y=-3 y=5/3 so the answer is not A

OpenStudy (anonymous):

ohh i understand now

OpenStudy (anonymous):

now lets plug in x=1 y - 2x = -1 x + 3y = 4 y-2(1)=-1 1+3y=4 y-2=-1 3y=3 y=1 y=1 so B is correct

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

:)